# Detuning Properties of RF Phase Modulation in an Electron Storage Ring

### v3

@@ II Theoretical model
In reference [2], they discuss the motion of the electron under the influence of RF phase modulation with a modulation frequency that is double the synchrotron oscillation frequency. In this section, we also discuss the motion of the electron in the same manner as [2], but by adding a detuning condition to the modulation frequency. We formulate the equation of motion in the case where the modulation frequency is slightly different from double the synchrotron oscillation frequency, and we derive the dependence of the electron’s amplitude and the bunch length on the detuning condition. First, we discuss the electron’s motion by solving its equation in the longitudinal direction using a linear approximation. In this approximation, the electron is affected by a linear restoring force that depends on time due to phase modulation. Next, we extend the discussion to a nonlinear approximation, in which the electron is affected by a nonlinear restoring force due to the large amplitude of the synchrotron oscillation. From these discussions, we derive a formula showing the dependence of bunch elongation on the difference in frequency between the double synchrotron frequency and the modulation frequency.
@@
Here $U$ is the synchrotron radiation loss, and $E$ is the beam energy. The phase $\phi_{m}$ represents the phase modulation term, and we now assume $\phi_{m}$ has the following modulation characteristics:
@@
where $\sigma_{\tau}$ is the natural bunch length. The derivation from Eq. (13) to Eq. (14) and (15) is summarized in Appendix A. From Eq. (15), the detuning frequency $\Delta\omega_{\rho}$ where the bunch length $\sigma_{\text{mod}}$ is equal to $\rho\sigma_{\tau}$ $(\rho>1)$ is given as
@@
$F_{\pm}(t)=\left(1+\alpha\left(1\pm\frac{2\alpha}{\Delta\omega}\right)t\right)^{2}.$ (19)
The derivation from Eq. (13) to Eq. (18) and (19) is summarized in Appendix A. The bunch length $\sigma_{\text{mod}}$ under the near resonance is given as
$\sigma_{\text{mod}}=\sigma_{\tau}\sqrt{1+\frac{1}{2}\left(\left(F_{+}(t)-1\right)e^{-\frac{2\alpha}{\Delta\omega}}+\left(F_{-}(t)-1\right)e^{\frac{2\alpha}{\Delta\omega}}\right)}.$ (20)
This means that the oscillation amplitude of the electron increases monotonically, and the bunch length diverges under the near resonance condition with the linear approximation.
To consider the motion of the electron under phase modulation precisely, we introduce nonlinear approximation by expanding the cosine part of the equation of motion shown in Eq. (2) up to the third-order terms, namely
@@
As a measure, we estimate the frequency width of the detuning curve in the KARA case. If we assume that the total accelerating voltage $V_{c}$, the radiation loss per turn $U_{0}$, and the synchrotron angular frequency $\omega_{s}$ are 1.36 MV, 622.4 keV, and $2\pi\times$34.3 kHz, which are the values we have evaluated from the experiment mentioned in section III. Regarding the phase modulation amplitude, 10.25 degrees was one of the modulation amplitude values in the experiments. As mentioned in section III, we activated phase modulation at one of the two RF stations. In this case, the parameter $\phi_{m0}$ in the model must be half of the set experimental value. That is, $\phi_{m0}$ must correspond to $\frac{1}{2}\times 10.25=5.13$ degrees. From these parameters, we evaluate the frequency width $\Delta f_{\sqrt{2}}=0.74$ kHz from Eq. (30). Concerning the negative frequency shift due to nonlinearity, we evaluate the negative detuning frequency $\delta f_{\sqrt{2}}=-0.38$ kHz from Eq. (31).
The Equation (27) also shows the relationship between the modulated bunch length and the frequency detuning, namely the detuning character. We write the phases occupied by the natural bunch length $\sigma_{\tau}$ and by the modulated bunch length $\sigma_{\text{mod}}$ as $\psi_{\tau}=\omega\sigma_{\tau}$ and $\psi_{\text{mod}}=\omega\sigma_{\text{mod}}$. For simplicity, assuming that the phase occupied by the modulated bunch length is small enough and approximate as $1-\left(\frac{\psi_{\text{mod}}}{2}\right)^{2}\approx 1$, Eq. (27) can be written as the following cubic equation for $\psi^{2}_{\text{mod}}$:
@@ III.1 Modulation frequency and amplitude
We have measured the bunch lengthening while changing the modulation frequency and amplitude of the phase modulation. Figs. 3 (a)-(c) show the streak camera images measured with a modulation amplitude of 10.25 degrees and a beam current of 88 mA. Fig. 3 (a) shows the streak camera image for the modulation frequency that causes the maximum bunch lengthening. Figs. 3 (b) and 3 (c) show the images with -1 kHz and +1 kHz detuning from the modulation frequency in Fig. 3 (a). The modulation frequency change of ±1 kHz from the maximum elongation condition in Fig. 3(a) clearly reduces the effect of bunch lengthening. As seen in these figures, the electron bunch length oscillates over time and the oscillation period is almost the same as half the synchrotron oscillation period $(\sim 15\ \mu\text{s})$.
(a) The maximum bunch lengthening condition
(b) -1 kHz detune from (a)
(c) +1 kHz detune from (a)
Figure 3: The experimental results of the streak camera measurement which show the change in the longitudinal beam profile. Figure 3 (a) (upper) shows the result at the maximum bunch lengthening condition with a modulation frequency of 67.0 kHz. Figures 3 (b) and (c) (middle and lower) show the streak camera results at modulation frequencies of 66.0 and 68.0 kHz, which correspond to the frequency conditions of -1 and +1 kHz from the maximum bunch lengthening condition. The beam current and the modulation amplitude are 88 mA and 10.25 degrees.
@@ III.3 Beam lifetime
Figure 7: The measured beam lifetime and bunch length with the modulation amplitudes of ((a), upper) 10.25 and ((b), lower) 16.42 degrees at the same beam current of 107 mA.
@@ IV Simulation
To consider the beam loading effect, we have treated both the phase advance and the amplitude of the beam-induced voltage $\vec{V_{b}}$ between adjacent bunches in multi-bunch filling conditions using the same method as in [2]. Because of the characteristics of the beam loading effect, initially, $\vec{V_{b}}$ inside each cavity can change with the beam revolution step. Therefore, we have performed enough revolution steps in the simulation to achieve a quasi-stable condition in which $\vec{V_{b}}$ and $\vec{V_{c}}$ can oscillate slightly but do not change drastically. Typically, we have performed the revolution step over a period of up to three times the radiation damping time. To get a realistic solution for the motion of the electron, we have introduced the radiation excitation into Eqs. (1) and (2) as a random walk of the energy spread, which has a Gaussian distribution with a standard deviation $\sigma_{r}$ as
@@ IV.1 Modulation frequency and amplitude
(a) The maximum bunch lengthening condition
(b) -1 kHz detune from (a)
(c) +1 kHz detune from (a)
Figure 9: The simulation results of the change in the longitudinal electron distribution in one bunch when the modulation frequency is set to maximize the bunch lengthening ($12002f_{s}-$ 1200 Hz. upper, Fig. 9 (a)), -1 kHz, and +1 kHz detuning from condition (a) (middle, Fig. 9 (b), and lower, Fig. 9 (c)). The 90 mA total beam current is set in the simulation, and all the other machine parameters are the same as in Fig. 8. The vertical scale measures the bunch length. The positive direction on the vertical axis corresponds to the head of the bunch. The blue and yellow colors indicate low and high electron density in the charge distribution, respectively.
@@ Appendix A Derivation of detuning character
$F_{\pm}(t)=\biggl{(}1+\alpha\Bigl{(}1\pm\frac{2\alpha}{\Delta\omega}\Bigr{)}t\biggr{)}^{2},$ (70)

### v4

# thanks: Present address: SLAC National Laboratory

@@ II Theoretical model
In reference [2], they discuss the motion of the electron under the influence of RF phase modulation with a modulation frequency that is double the synchrotron oscillation frequency. In this section, we also discuss the motion of the electron in the same manner as [2], but by adding a detuning condition to the modulation frequency [3]. We formulate the equation of motion in the case where the modulation frequency is slightly different from double the synchrotron oscillation frequency, and we derive the dependence of the electron’s amplitude and the bunch length on the detuning condition. First, we discuss the electron’s motion by solving its equation in the longitudinal direction using a linear approximation. In this approximation, the electron is affected by a linear restoring force that depends on time due to phase modulation. Next, we extend the discussion to a nonlinear approximation, in which the electron is affected by a nonlinear restoring force due to the large amplitude of the synchrotron oscillation. From these discussions, we derive a formula showing the dependence of bunch elongation on the difference in frequency between the double synchrotron frequency and the modulation frequency.
@@
Here $U$ is the synchrotron radiation loss, and $E$ is the beam energy. The phase $\phi_{m}$ represents the phase modulation term, and we now assume $\phi_{m}$ has the following modulation characteristics [2, 3]:
@@
where $\sigma_{\tau}$ is the natural bunch length. The derivation from Eq. (13) to (15) is summarized in Appendix A. From Eq. (15), the detuning frequency $\Delta\omega_{\rho}$ where the bunch length $\sigma_{\text{mod}}$ is equal to $\rho\sigma_{\tau}$ $(\rho>1)$ is given as
@@
$F_{+}(t)=\left(1\pm\alpha\left(1+\frac{2\alpha}{\Delta\omega}\right)t\right)^{2},$ (19)
$F_{-}(t)=\left(1\pm\alpha\left(1-\frac{2\alpha}{\Delta\omega}\right)t\right)^{2}.$
The derivation from Eq. (13) to (18) is summarized in Appendix A. This means that the oscillation amplitude of the electron increases monotonically, and the bunch length diverges under the near resonance condition with the linear approximation.
To consider the motion of the electron under phase modulation precisely, we introduce nonlinear approximation by expanding the cosine part of the equation of motion shown in Eq. (2) up to the third-order terms, namely [3]
@@
As a measure, we estimate the frequency width of the detuning curve in the KARA case. Now, we assume that the total accelerating voltage $V_{c}$, the radiation loss per turn $U_{0}$, and the synchrotron angular frequency $\omega_{s}$ are 1.36 MV, 622.4 keV, and $2\pi\times$34.3 kHz, which are the values we have evaluated from the experiment mentioned in section III. Regarding the phase modulation amplitude, 10.25 degrees was one of the modulation amplitude values in the experiments. As mentioned in section III, we activated phase modulation at one of the two RF stations. In this case, the parameter $\phi_{m0}$ in the model must be half of the set experimental value. That is, $\phi_{m0}$ must correspond to $\frac{1}{2}\times 10.25=5.13$ degrees. From these parameters, we evaluate the frequency width $\Delta f_{\sqrt{2}}=0.74$ kHz from Eq. (29). Concerning the negative frequency shift due to nonlinearity, we evaluate the negative detuning frequency $\delta f_{\sqrt{2}}=-0.38$ kHz from Eq. (30).
The Equation (26) also shows the relationship between the modulated bunch length and the frequency detuning, namely the detuning character. We write the phases occupied by the natural bunch length $\sigma_{\tau}$ and by the modulated bunch length $\sigma_{\text{mod}}$ as $\psi_{\tau}=\omega\sigma_{\tau}$ and $\psi_{\text{mod}}=\omega\sigma_{\text{mod}}$. For simplicity, assuming that the phase occupied by the modulated bunch length is small enough to approximate as $1-\left(\frac{\psi_{\text{mod}}}{2}\right)^{2}\approx 1$, Eq. (26) can be written as the following cubic equation for $\psi^{2}_{\text{mod}}$:
@@ III.1 Modulation frequency and amplitude
We have measured the bunch lengthening while changing the modulation frequency and amplitude of the phase modulation. Figures 3(a)-3(c) show the streak camera images measured with a modulation amplitude of 10.25 degrees and a beam current of 88 mA. Figure 3(a) shows the streak camera image for the modulation frequency that causes the maximum bunch lengthening. Figures 3(b) and 3(c) show the images with -1 kHz and +1 kHz detuning from the modulation frequency in Fig. 3(a). The modulation frequency change of $\pm$1 kHz from the maximum elongation condition in Fig. 3(a) clearly reduces the effect of bunch lengthening. As seen in these figures, the electron bunch length oscillates over time and the oscillation period is almost the same as half the synchrotron oscillation period $(\sim 15\ \mu\text{s})$.
Figure 3: The experimental results of the streak camera measurement, which show the change in the longitudinal beam profile. (a) shows the result at the maximum bunch lengthening condition with a modulation frequency of 67.0 kHz. (b), (c) show the streak camera results at modulation frequencies of 66.0 and 68.0 kHz, which correspond to the frequency conditions of -1 and +1 kHz from the maximum bunch lengthening condition. The beam current and the modulation amplitude are 88 mA and 10.25 degrees.
@@ III.3 Beam lifetime
Figure 7: The measured beam lifetime and bunch length with the modulation amplitudes of (a) 10.25 and (b) 16.42 degrees at the same beam current of 107 mA.
@@ IV Simulation
To consider the beam loading effect, we have treated both the phase advance and the amplitude of the beam-induced voltage $\vec{V_{b}}$ between adjacent bunches in multi-bunch filling conditions using the same method as in [2]. Because of the characteristics of the beam loading effect, initially, $\vec{V_{b}}$ inside each cavity can change with the beam revolution step. Therefore, we have performed enough revolution steps in the simulation to achieve a quasi-stable condition in which $\vec{V_{b}}$ and $\vec{V_{c}}$ can oscillate slightly but do not change drastically. Typically, we have performed the revolution step over a period of up to three times the radiation damping time. To get a realistic solution for the motion of the electron, we have introduced the radiation excitation into Eqs. (1) and (2) as a random walk of the energy spread, which has a Gaussian distribution with a standard deviation $\sigma_{r}$ as [2]
@@ IV.1 Modulation frequency and amplitude
Figure 9: The simulation results of the change in the longitudinal electron distribution in one bunch when the modulation frequency is set to maximize the bunch lengthening [$12002f_{s}-$ 1200 Hz;(a)], -1 kHz, and +1 kHz detuning from condition (a) [(b), (c)]. The 90 mA total beam current is set in the simulation, and all the other machine parameters are the same as in Fig. 8. The vertical scale measures the bunch length. The positive direction on the vertical axis corresponds to the head of the bunch. The blue and yellow colors indicate low and high electron density in the charge distribution, respectively.
@@ Appendix A Derivation of detuning character
$F_{+}(t)$ $=\biggl{(}1\pm\alpha\Bigl{(}1+\frac{2\alpha}{\Delta\omega}\Bigr{)}t\biggr{)}^{2},$ (69)
$F_{-}(t)$ $=\biggl{(}1\pm\alpha\Bigl{(}1-\frac{2\alpha}{\Delta\omega}\Bigr{)}t\biggr{)}^{2},$